I am really having a hard time trying to figure out this problem... √ -x+7 = x-1
Answer by edjones(8007)

sqrt(-x+7)=x-1 I assume this is the question.
-x+7=(x-1)^2 Square each side.
-x+7=x^2-2x+1
x^2-2x+1+x-7=0
x^2-x-6=0
(x-3)(x+2)=0
x=3, x=-2
.
Check:
sqrt(-3+7)=3-1
sqrt(4)=2 true
.
sqrt(2+7)=-2-1
sqrt(9)=+-3 true
.
Ed